how do i solve 3/4+5-7
Answer by Stxfball(4)

Well you take into account order of operations PEMDAS Parenthasese Exponent Multiplication(interchangeable withone another)Division Addition(interchangeable withone another)Subtraction. so 3/4+5-7 3/4=.75 so now it is .75+5-7= 5.75-7 and from there the final answer -1.25